# Correlations of Feed-down Hadrons in a Thermal Model

### v1

@@ I Introduction
The goal of this work is to estimate the impact of decay FD onto measurable “stable particles” based on a simple version of the thermal hadron gas model Andronic et al. (2009). Our discussion is framed in the context of the grand canonical model (GCM) and capitalizes on existing tools developed towards the estimation of the role of FD on single particle production Chojnacki et al. (2012). It is assumed that hot QCD matter formed in A–A collisions cools down sufficiently slowly, as the collision system expands, to remain in thermal equilibrium at all times until chemical and kinetic freeze out. It is additionally assumed that all hadrons, irrespective of their mass and flavor content, reach chemical freeze-out at the same proper time, so one can neglect the fact that heavier particles stop being formed earlier (by detailed balance) than their lighter cousins.
@@ III Results
Focusing on the impact of FD at $T=$ 160 MeV, which corresponds to typical freeze-out temperatures obtained by blast wave model fits of single species momentum spectra at both RHIC (top energy) and LHC Xu and Ko (2017), one notes that FD contributions increase proton density by a factor of $\sim 3.1$. This is relevant to note given that measurement of net-baryon yield fluctuations, particularly net-protons, are used as a proxy for net-baryon fluctuations aimed at a determination of chemical susceptibilities and searches for the critical point of nuclear matter. Although, strong decays preserve the net baryon number, they do not preserve the net proton number: baryon decays feature p, n, or $\Lambda$ as end points (in the context of this discussion). The density of observable protons $\rho_{\rm p}^{\rm TH+FD}$, assuming perfect efficiency and acceptance, is thus related to that of the total primary baryon density $\rho_{\rm B}^{\rm TH}$ according to
@@
Correlations involving one proton are arguably the simplest. No resonant state decay prominently into baryon anti-baryon pairs. However, the annihilation of $\rm c\bar{c}$ states, such as $\eta_{\rm c}$, $\rm J/\psi$, etc, do decay into baryon anti-baryon channels plus other particles, albeit with branching fractions of order $10^{-3}$ or smaller. The thermal production of $\rm c\bar{c}$ states is at best extremely weak given their large mass. Their decays thus contribute negligibly to the baryon yield and fluctuations of the net baryon number on an event-by-event basis. Consequently, as shown in the bottom panel of Fig. 3, proton from decays are most often associated to pions, with the strength of correlations being the largest for neutral pions $\pi^{0}$ followed closely by that of $\pi^{+}$ and $\pi^{-}$. Again, one notes that while (strong) decays do conserve quantum numbers (and flavor), they do not necessarily yield charge and baryon balancing pairs. We also observe that similar effects play out for correlations involving $\Lambda$ baryons.
@@
Within the context of processes driven by the strong interaction, conservation laws dictate that the baryon number, the charge and more generally the flavor content of produced hadrons must be balanced. However, charge, flavor, and baryon balance are not required to be realized by the correlated emission of only two particles. It is then possible for the emission of a proton to be charge balanced by the production of $\pi^{-}$ even though this does not satisfy the baryon balance requirement. It is indeed possible to balance the charge, the baryon number, and quark flavor independently. This leads to simple sum rules of the form $\sum_{\alpha}B(\alpha|\beta)=1$, where $B(\alpha|\beta)$ is the integrated balance function of species $\beta$ being balanced by the emission of species $\alpha$ Pruneau et al. (2023, 2024). The production of a hadron $\beta$ can be charge balanced (or alternatively baryon balanced, etc) by several distinct hadrons and the sum of the integral of the balance functions must add to unity. Such simple sum rules are evidently violated in the context of the GCE given (a) the emission of thermal hadrons are uncorrelated and (b) strong decays induces correlations that do not balance charge, baryon number, or quark flavors (although they do individually conserve quantum numbers). It is thus of interest to explicitly examine the impact of decays on balance functions by normalizing the 2-cumulants $C_{2}(\alpha|\beta)$ by their sum $\sum_{\alpha}C_{2}(\alpha|\beta)$, as illustrated in Fig. 5.
@@ Acknowledgments
The authors acknowledge the use of computer code and hadron data from the Therminator2 model and discussions with the creators of this model. The authors are also thankful for discussions with S. Voloshin. This work was supported in part by grant No. DE-FG02-92ER40713, No. DE-SC0021969, grant No. SR/MF/PS-02/2021-IITB (E-37126), and grant No. PN-III-P4-PCE-2021-0390. C.S. acknowledges a DOE Office of Science Early Career Award.

### v2

@@ I Introduction
The goal of this work is to estimate the impact of decay FD onto measurable “stable particles” based on a simple version of the thermal hadron gas model Andronic et al. (2009). Our discussion is framed in the context of the grand canonical model (GCM) and capitalizes on existing tools developed towards the estimation of the role of FD on single particle production Chojnacki et al. (2012). It is assumed that hot QCD matter formed in A–A collisions cools down sufficiently slowly, as the collision system expands, to remain in thermal equilibrium at all times until chemical and kinetic freeze out. It is additionally assumed that all hadrons, irrespective of their mass and flavor content, reach chemical freeze-out at the same proper time, so one can neglect the fact that heavier particles stop being formed earlier (by detailed balance) than their lighter cousins. Effects of quantum number conservation have been addressed by other authors and are not considered in this work Vovchenko et al. (2020); Braun-Munzinger et al. (2024).
@@ III Results
Focusing on the impact of FD at $T=$ 160 MeV, which corresponds to typical freeze-out temperatures obtained by blast wave model fits of single species momentum spectra at both RHIC (top energy) and LHC Xu and Ko (2017), one notes that FD contributions increase proton density by a factor of $\sim 3.1$. This is relevant to note given that measurement of net-baryon yield fluctuations, particularly net-protons, are used as a proxy for net-baryon fluctuations aimed at a determination of chemical susceptibilities and searches for the critical point of nuclear matter. Although, strong and weak decays preserve the net baryon number, they do not preserve the net proton number: baryon decays feature p, n, or $\Lambda$ as end points (in the context of this discussion). The density of observable protons $\rho_{\rm p}^{\rm TH+FD}$, assuming perfect efficiency and acceptance, is thus related to that of the total primary baryon density $\rho_{\rm B}^{\rm TH}$ according to
@@
Correlations involving one proton are arguably the simplest. No resonant state decay prominently into baryon anti-baryon pairs. However, the annihilation of $\rm c\bar{c}$ states, such as $\eta_{\rm c}$, $\rm J/\psi$, etc, do decay into baryon anti-baryon channels plus other particles, albeit with branching fractions of order $10^{-3}$ or smaller. The thermal production of $\rm c\bar{c}$ states is at best extremely weak given their large mass. Their decays thus contribute negligibly to the baryon yield and fluctuations of the net baryon number on an event-by-event basis. Consequently, as shown in the bottom panel of Fig. 3, proton from decays are most often associated to pions, with the strength of correlations being the largest for neutral pions $\pi^{0}$ followed closely by that of $\pi^{+}$ and $\pi^{-}$. Again, one notes that while (E&M, weak and strong) decays do conserve baryon numbers, they do not necessarily yield charge and baryon balancing pairs. We also observe that similar effects play out for correlations involving $\Lambda$ baryons.
@@
Within the context of processes driven by the strong interaction, conservation laws dictate that the baryon number, the charge and more generally the flavor content of produced hadrons must be balanced. However, charge, flavor, and baryon balance are not required to be realized by the correlated emission of only two particles. It is then possible for the emission of a proton to be charge balanced by the production of $\pi^{-}$ even though this does not satisfy the baryon balance requirement. It is indeed possible to balance the charge, the baryon number, and quark flavor independently. This leads to simple sum rules of the form $\sum_{\alpha}B(\alpha|\beta)=1$, where $B(\alpha|\beta)$ is the integrated balance function of species $\beta$ being balanced by the emission of species $\alpha$ Pruneau et al. (2023, 2024). The production of a hadron $\beta$ can be charge balanced (or alternatively baryon balanced, etc) by several distinct hadrons and the sum of the integral of the balance functions must add to unity. Such simple sum rules are evidently violated in the context of the GCE given (a) the emission of thermal hadrons are uncorrelated and (b) decays induces correlations that do not balance charge, baryon number, or quark flavors (although they do individually conserve quantum numbers). It is thus of interest to explicitly examine the impact of decays on balance functions by normalizing the 2-cumulants $C_{2}(\alpha|\beta)$ by their sum $\sum_{\alpha}C_{2}(\alpha|\beta)$, as illustrated in Fig. 5.
@@ Acknowledgments
The authors acknowledge the use of computer code and hadron data from the Therminator2 model and discussions with the creators of this model. The authors are also thankful for discussions with S. Voloshin and V. Vovchenko. This work was supported in part by grant No. DE-FG02-92ER40713, No. DE-SC0021969, grant No. SR/MF/PS-02/2021-IITB (E-37126), and grant No. PN-III-P4-PCE-2021-0390. C.S. acknowledges a DOE Office of Science Early Career Award.